If |−2X + 9| = 3 then the possible value of |−X| −X^2  would be:


(A) 30 (B) -30 (C) -42 (D) 42

Grade:Upto college level

2 Answers

Swetha
10 years ago
as |-2X+9| = 3, -2X+9 = 3 or -2X+9 = -3. first let -2X+9 = 3 ==> -2X = -6 ==> X = 3. so |-X| -X^2 = -6. now let -2X+9 = -3. ==> X = -6 . so |-X| -X^2 = -30. hence the answer is option B
akshita
3 years ago
 

Given  is |−2X+9|=3

removing mod operator with both sign
case 1 : if |−2X+9|>=0 mod will open with positive sign  i.e. (-2x+9)>=0 

                   hence x

              ⇒ (−2X+9)=3

              ⇒  x=3  It is rejected as x should be less than 4.5 to satisfy the condition for case 1 

 

case 2 : if |−2X+9|

              hence x>4.5

              ⇒     (−2X+9)=-3   

              ⇒     x=6 This value is accepted as it satisfies the condition x>4.5 of case 2

now put value x=6 in 

|-X|-X^2 = |-6|-36= -30

 


 Thus B is the correct option